OpenStudy (anonymous):

In the scene in an action movie, a stuntman jumps from the top of one building to the top of another building 4.0 m away. After a running start, he leaps at an angle of 15 degrees with respect to the flat roof while traveling at a speed of 5.0 m/s. Will he make it to the other roof, which is 2.5 m shorter than the building he jumps from. I've been stuck on it for so long...I can't find out what (a) in the equation would be. Please help.

OpenStudy (anonymous):

find r \[r=\frac{ V _{0}^{2}\sin 2\theta }{g }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Y : 0 =2.5 + 5sin15*t - (1/2)*g*t^2 solve for t; use in: X: x= 5cos15*t if x>= 4, he makes it.

OpenStudy (anonymous):

Yes! Thank you!!

OpenStudy (anonymous):

sure:)

OpenStudy (anonymous):

I might need more help tonight if you're on it would be amazing if you could help me. I also am having problems with relative velocity